eviant:

PLS HELP!!! Let f(x) = x2 and g(x) = x − 3. Evaluate (g ∘ f)(−2).

TheSmartOne:

So first find f(-2) and then plug that number into g(x)

sillybilly123:

OR.....just do the composition

eviant:

@dude

eviant:

@Shadow

TheSmartOne:

You're right, Billy. In this case, it would be just as easy (and not messy at all) to find the composition.

TheSmartOne:

Also, eviant, if you're not going to respond and participate in your own question after receiving help and instead tag other users, then you're mostly likely not going to receive help from others. QuestionCove is not a resource to give you answers but rather to help you learn. We can teach you how to answer this simple question so that when you're assignment has 10 of these questions, you'll get them all right instead of having to post 10 questions.

eviant:

@TheSmartOne how do I find f(x)=X^2?

TheSmartOne:

(g ∘ f)(−2) = g(f(-2)) So we want to do the innermost parenthesis f(-2) = ?? we know that f(x) = x^2 so to find f(-2) we just plug -2 in the place of x and solve so what is (-2)^2 = ??

eviant:

-4

TheSmartOne:

Uh no

eviant:

aren't you saying to square -2?

eviant:

sorry, I meant 4

TheSmartOne:

Yes, that's correct so we go from \(g( \color{red}{f(-2)}) = g(\color{red}{4})\) Do you understand so far?

eviant:

yes, I think

eviant:

so, I sub -2 for f?

TheSmartOne:

So when we did f(-2) that was (-2)^2 now we need to find g(4) g(x) = x - 3 g(4) = ?? just replace the x with 4 and solve

eviant:

g(x)=4-3 g(4)=1

eviant:

is that correct?

TheSmartOne:

Correct! And that's your answer

eviant:

thank you so much, also I apologie for not being here first, next time I'll try to post my question when I have time to answer it

eviant:

apologize*

sillybilly123:

@TheSmartOne Doing both is the Counsel of Perfection :)

TheSmartOne:

Definitely :)

TheSmartOne:

And no problem, eviant! Hopefully you learned something :)